OpenStudy (anonymous):

8. Determine the length of the leg of a 45* – 45* – 90* triangle with a hypotenuse length of 15 inches. WILL FAN AND MEDAL!

OpenStudy (anonymous):

what is the asterix for?

OpenStudy (anonymous):

jkjk its for the degrees

OpenStudy (anonymous):

45^2+45^2=90^2

OpenStudy (anonymous):

Ok, so the degrees are being subtracted?

OpenStudy (anonymous):

no thats just used as a comma

OpenStudy (anonymous):

if i know the lengths of the sides of a 45-45 -90 right triangle, then i can find the hypotenuse by multiplying by \[\sqrt{2}\]

OpenStudy (anonymous):

if i know the length of the hypotenuse, i can find the length of the sides by dividing by the same

OpenStudy (anonymous):

\[15/\sqrt{2}\]

OpenStudy (campbell_st):

use what you know... the triangle is isoceles with 2 equal sides  just solve for a

OpenStudy (anonymous):

it is just 15/sqrt2

OpenStudy (anonymous):

@babygurl47672

OpenStudy (anonymous):

@pecbotad